a gondola travels 3250 feet to reach the middle station. the middle station is 1100 feet higher than the base station. the suspension railway then travels another 920 feet to the summit station. how high is the summit station above the base station?
Answer by addingup(3677)

The rise is 1100 and the run 3250. And since it's rise/run:
It's rising 1100/3250= 0.338 feet for every foot of run. Therefore, an additional 920 feet of run would increase the rise by:
920*0.338= 310.96
The summit station is:
1100+310.96= 1410.96 feet above the base (you can round it to 1411 feet if they allow you to do that)
J